OpenStudy (vshiroky):

Find a ratio equivalent to each ratio. Then use the ratios to write a proportion. 7/9 11/12 I don't understand

OpenStudy (vshiroky):

I think the first can be 14/18 which is 7:9::14:18

OpenStudy (anonymous):

\[\frac{7}{9}=\frac{14}{18}=\frac{21}{27}=\frac{28}{36}=...\]

OpenStudy (vshiroky):

do I need to write it how I wrote it?

OpenStudy (vshiroky):

I dont understand what it means for proportion

OpenStudy (anonymous):

i do not understand the part that says Then use the ratios to write a proportion.

OpenStudy (anonymous):

maybe it means for example \[\frac{7}{9}=\frac{14}{18}\\ 7:9::14:19\]

OpenStudy (anonymous):

typo there

OpenStudy (anonymous):

\[\frac{7}{9}=\frac{14}{18}\\ 7:9::14:18\]

OpenStudy (anonymous):

that is my best guess, i would not bet $16 on it, only $10

OpenStudy (vshiroky):

Lol ok thank you